# End-to-End Evaluation for Low-Latency Simultaneous Speech Translation

### v3

@@ 1 Introduction
  • A framework [^2] https://gitlab.kit.edu/kit/isl-ai4lt/lt-middleware/ltpipeline for low-latency speech translation with dynamic latency adjustment
@@ 5 Evaluation Framework
We evaluate our system in an end-to-end fashion. That is, given an input audio, we send it to the system and evaluate the final returned transcript and translation. We provide an evaluation framework [^4] https://gitlab.kit.edu/kit/isl-ai4lt/lt-middleware/lt-evaluation that assess the system in different aspects and logs the results to categorized experiments on an UI board using MLflow Zaharia et al. (2018). We consider different evaluation metrics as follows.
@@ 6.1 Evaluation data
Test data Lang. pair Hours # Utterances
@@ 7.2 Revision mode vs fixed mode
Second, we report the results of comparing the revision mode to the fixed mode with different $LA2\_chunk\_size$ values when performing cascaded translation on the en-de ACL dev set. As can be seen in Figure 5, in general, revision mode has better BLEU score yet worse latency than fixed mode. This is expected, since for the revision mode, when more input audio is available, the system can correct its previous output, thus ending up having better translation quality yet worse latency due to the additional re-translation overhead.

### v4

@@ 1 Introduction
  • A framework [^2] https://git.scc.kit.edu/isl/lt-middleware/ltpipeline for low-latency speech translation with dynamic latency adjustment
@@ 5 Evaluation Framework
We evaluate our system in an end-to-end fashion. That is, given an input audio, we send it to the system and evaluate the final returned transcript and translation. We provide an evaluation framework [^4] https://git.scc.kit.edu/isl/lt-middleware/lt-evaluation that assess the system in different aspects and logs the results to categorized experiments on an UI board using MLflow Zaharia et al. (2018). We consider different evaluation metrics as follows.
@@ 6.1 Evaluation data
Test data Lang. pair Hours # Utt.
@@ 7.2 Revision mode vs fixed mode
Second, we report the results of comparing the revision mode to the fixed mode with different $LA2\_chunk\_size$ values when performing cascaded translation on the en-de ACL dev set. As can be seen in Figure 5, in general, revision mode has better BLEU score and better latency than fixed mode. The better BLEU score comes from the revision mode being able to correct its previous output when more audio input is provided as time goes by.